what is the answer to 2x-y-z=15, 4x+5y+2z=10, -x-4y+3z=-20
Answer by Alan3354(69443)

2x-y-z=15
4x+5y+2z=10
-x-4y+3z=-20
-------------
x = 5
y = 0
z = -5